# Re-boosting Self-Collaboration Parallel Prompt GAN for Unsupervised Image Restoration

### v1

@@ 4.2 Implementation Details
To optimize the proposed network, we use the Adam optimizer with $\beta_{1}$=0.9, $\beta_{2}$=0.999, and an initial learning rate of $2\times 10^{-4}$. The proposed models are implemented using PyTorch and trained on two Nvidia GeForce RTX 3090 GPUs. For the denoising task, the batch size and patch size are set to 6 and 112, respectively. We set $\lambda_{\text{BGM}}$ in Eq. 4 to 6 and $\lambda_{\text{SSIM}}$ in Eq. 5 to 1. We use a ResNet with 6 residual blocks as the generator, a PatchGAN [47] as the discriminator, and a Restormer [6] as the restorer. For the deraining task, following prior work [21], we set the batch size to 1 and the patch size to 256. The NAFNet [5] as our restorer.
@@ 4.3 Image Denoising
TABLE III: Deraining results of several competitive methods on Rain100L and RealRainL. The highest PSNR(dB)/SSIM is highlighted in red, while the second is in blue.
@@
Restormer [6] No/CVPR 2022 38.34 0.9830 40.90 0.9850 -- --
@@
Figure 11: Restoration results of our method against other competing methods on the RealRainL [26] dataset.
@@
TABLE IV: Ablation studies on the proposed modules. V1: (U) Conventional GAN-based unsupervised denoising network only with unpaired synthesis; V2: V1 + BGMloss; V3: V1 + BGMloss + PL module; V4: (S) SGabNet (V1 + BGMloss + PL module + self-synthesis) V5: (P) P2GAN (our baseline).
@@
TABLE V: Ablation studies of Re-boosting module on SC strategy (Reb-SC) on SIDD Validation, SIDD Benchmark, Rain100L, and RealRainL. The N is the number of augmentation images.
@@ 4.4 Image Deraining
Quantitative Comparison. Quantitative Comparison. We evaluate the deraining performance of our method on the Rain100L [24], RealRainL [26] and Rain12 [25] datasets. We note that there are few unsupervised deraining methods with source codes for performance evaluation. Table III shows that RSCP2GAN outperforms all unsupervised approaches. Compared to DerainCycleGAN [20] and DCDGAN [21], our method achieves a PSNR gain of 2.52 dB and 2.19 dB, and an SSIM gain of 0.024 and 0.019 on the Rain100L test set. Our method also performs comparably to some supervised methods like SPA-Net [72] and DDN [70].
Qualitative Comparison. Fig. 11 visually compares deraining methods on the Rain100L and RealRainL datasets. RSCP2GAN achieves favorable deraining results, effectively removing rain streaks where other methods fail and avoiding chromatic aberration seen with DCD-Net [21]. On the RealRainL dataset, CycleGAN tends to oversharpen image details, and DCD-GAN causes severe blurriness. The second row of images confirms that RSCP2GAN provides the best visual results for image deraining.
@@ 4.5 Ablation Study
Effectiveness of the Proposed Framework. We validate the effectiveness of the P2GAN structure as described in Table IV for the denoising task. Here, V1 represents a GAN-based unsupervised denoising network with only unpaired synthesis. In addition, V2 extends V1 by adding the BGM loss, while V3 further includes the PL module. V4 introduces the branch U-S, and V5 is our baseline (P2GAN).
Table IV shows that adding the BGM loss to the GAN-based unsupervised network (V1 to V2) results in a 0.17 dB PSNR improvement. Incorporating the PL module (V2 to V3) yields an additional PSNR increase of approximately 0.23 dB, highlighting the effectiveness of the PL module in enhancing synthetic image quality. When adding the “self-synthesis” constraint to V3 to obtain V4, there is a notable improvement of 0.45 dB PSNR on the SIDD Benchmark, indicating that combining “self-synthesis” with unpaired synthesis enhances network training and restorer performance. Comparing P2GAN with V4, we observe a substantial performance boost in P2GAN, with PSNR gains of 0.53 dB on the SIDD Validation. This demonstrates that P2GAN produces more realistic synthetic degraded images, improving the restorer’s performance.
@@
Effectiveness of the Number of Augmentation Images (N) in the Re-boosting (ReB) Module. Table V shows that Reb-SC improves performance across all datasets. For $N$=2, the improvement is approximately 0.2 dB; for $N$=4, it is about 0.3 dB. However, when N=8, the performance improvement is less significant than $N$=4.

### v2

@@ 4.1 Datasets
Desnowing Task. We train and test our model on commonly used desnowing datasets: CSD [27], and Snow100K [28]. The CSD contains 8000 image pairs for training and 2000 image pairs for testing. For the Snow100K, we use the presented 50000 image pairs for training and randomly select 2000 samples for testing. To ensure fair comparison under an unsupervised training setting, we apply the same training protocol as used in our denoising and deraining experiments.
@@ 4.2 Implementation Details
To optimize the proposed network, we use the Adam optimizer with $\beta_{1}$=0.9, $\beta_{2}$=0.999, and an initial learning rate of $2\times 10^{-4}$. The proposed models are implemented using PyTorch and trained on two Nvidia GeForce RTX 3090 GPUs. For the denoising task, the batch size and patch size are set to 6 and 112, respectively. We set $\lambda_{\text{BGM}}$ in Eq. 4 to 6 and $\lambda_{\text{SSIM}}$ in Eq. 6 to 1. We use a ResNet with 6 residual blocks as the generator, a PatchGAN [57] as the discriminator, and a Restormer [6] as the restorer. For the deraining and desnowing tasks, following prior work [21], we set the batch size to 1 and the patch size to 256. The NAFNet [5] as our restorer.
@@ 4.3 Image Denoising
TABLE III: Deraining results of several competitive methods on Rain100L, RealRainL, and Rain12.
@@
Restormer [6] No/CVPR 2022 37.57 0.9740 40.90 0.9850 -- --
PromptIR [51] NeurIPS 2024 38.34 0.9830 36.99 0.9730 35.09 0.9450
NeRD-Rain-S [86] CVPR 2024 42.00 0.9900 38.64 0.9790 35.39 0.9420
@@
DCDGAN [21] + ConvIR [87] Yes/TPAMI 2024 32.56 0.9547 31.23 0.9423 32.15 0.9301
DCDGAN [21] + NeRD [53] Yes/CVPR 2024 32.34 0.9606 31.41 0.9435 32.36 0.9297
@@
Figure 11: Restoration results of our model against other competing methods on the CSD [27] dataset.
@@
TABLE IV: Desnowing results of several competitive methods on CSD and Snow100K.
Methods GAN-based CSD Snow100K-S
[…]
Supervised MGF [88] No/ICNIP 2013 13.98 0.6700 24.32 0.7700
DesnowNet [28] No/TIP 2018 20.13 0.8100 32.33 0.9500
JSTASR [89] No/ECCV 2020 27.96 0.8800 31.40 0.9012
InvDSNet [90] No/TCSVT 2023 31.85 0.9600 34.39 --
LMQFormer [52] No/TCSVT 2023 32.64 0.9630 34.19 0.9436
PEUNet [54] No/TCSVT 2025 38.27 0.9900 36.76 --
ConvIR [87] No/TPAMI 2024 39.10 0.9900 36.37 0.9703
Unsupervised CycleGAN [57] Yes/CVPR 2020 16.72 0.6975 20.71 0.7639
DerainCycleGAN [20] Yes/TIP 2021 20.67 0.7963 22.79 0.8024
DGP-Cyc-GAN [66] Yes/ICPR 2022 21.35 0.8042 23.41 0.8072
DerainCycleGAN + ConvIR [87] Yes/TPAMI 2024 22.18 0.8132 23.97 0.8139
DerainCycleGAN + MWFormer [91] Yes/CVPR 2024 22.42 0.8179 24.45 0.8253
RSCP2GAN (Ours) Yes/-- 23.38 0.8462 25.97 0.8398
TABLE V: Ablation studies on the proposed modules. V1: (U) Conventional GAN-based unsupervised denoising network only with unpaired synthesis; V2: V1 + BGMloss; V3: V1 + BGMloss + PL module; V4: (S) SGabNet (V1 + BGMloss + PL module + self-synthesis) V5: (P) P2GAN (our baseline).
@@
TABLE VI: Ablation studies of Re-boosting module on SC strategy (Reb-SC) on SIDD Validation, SIDD Benchmark, Rain100L, and RealRainL. The N is the number of augmentation images.
@@ 4.4 Image Deraining
Quantitative Comparison. We evaluate the deraining performance of our method on the Rain100L [24], RealRainL [26] and Rain12 [25] datasets. We note that there are few unsupervised deraining methods with source codes for performance evaluation. Besides the existing unsupervised deraining frameworks, DerainingCycleGAN [20], NLCL [68], and DCDGAN [21], we add two new deraining networks, ConvIR [87] (TPAMI2024) and NeRD [53] (CVPR2024) as comparison methods. For fair unsupervised comparison, these methods are incorporated into the DCDGAN framework as the restorer, and are trained using the same datasets as ours. Table III shows that RSCP2GAN outperforms all unsupervised approaches. Compared to DerainCycleGAN [20] and DCDGAN [21], our method achieves a PSNR gain of 2.52 dB and 2.19 dB, and an SSIM gain of 0.024 and 0.019 on the Rain100L test set. Our method also performs comparably to some supervised methods like SPA-Net [84] and DDN [82]. Equipped with more advanced restorers, both DCDGAN+ConvIR and DCDGAN+NeRD achieve better performance than the original DCDGAN. However, our RSCP2GAN (using the traditional NAFNet as the restorer) still achieves the best results in three test sets. This indicates that the improvement from superior restorers is limited compared to advancements in restoration mechanisms, which further highlights the effectiveness of our SC and Reb-SC strategies.
Qualitative Comparison. Fig. 10 visually compares deraining methods on the Rain100L dataset. As shown, traditional methods such as CycleGAN and DCDGAN struggle to recover fine textures and complex degradations, often leaving visible artifacts, over-smoothed regions, and broken structures. The introduction of structural enhancement modules like ConvIR and NeRD leads to moderate improvements in the DCDGAN variants, resulting in clearer local details and slightly improved structural consistency. However, these methods still exhibit texture distortion and edge instability, particularly in scenes with dense textures or complex backgrounds. In contrast, RSCP2GAN consistently delivers sharper and more natural restoration results across all samples. It better preserves detailed textures, such as animal fur and background foliage, and maintains strong edge continuity, effectively suppressing artifacts and avoiding over-smoothing. In the zoomed-in regions, RSCP²GAN’s outputs closely resemble the ground truth (GT), demonstrating superior capability in both degradation modeling and restoration learning. These observations confirm the robustness and generalization ability of our framework in real-world restoration scenarios.
4.5 Image Desnowing
We evaluate the proposed method and state-of-the-art approaches on image desnowing benchmark datasets. The results of the supervised methods from previous works are presented as a reference. The comparison focuses on the performance of unsupervised methods.
Quantitative Comparison. Since there is limited exploration of unsupervised desnowing in existing work, we adopt the classic unsupervised restoration architecture, DerainCycleGAN, as a baseline for comparison, and replace its restoration module with the latest restoration networks, ConvIR [87] and MWFormer [91]. Unsupervised training consistent with the proposed method is conducted using the aforementioned desnowing dataset. In contrast, our framework still uses traditional NAFNet as the restorer, which is consistent with the deraining task. As shown in the Table IV, for the desnowing task, despite not using the latest restoration networks, our method still achieves a significant advantage on both test sets, thanks to the design of our RSCP2GAN framework. Specifically, compared to DerainCycleGAN [20], DGP-Cyc-GAN [66], and DerainCycleGAN + MWFormer [91], our model provides a PSNR gain of 2.71 dB, 2.03 dB, and 0.96 dB, and an SSIM gain of 0.050, 0.042, and 0.028 on the CSD test set.
Qualitative Comparison. The Fig. 11 presents a visual comparison between existing methods and our approach on the CSD dataset. Our method can effectively remove snowflakes, while others tend to leave residual snow. Additionally, existing methods often produce blurry results in snow removal scenarios with complex textures, whereas our approach preserves more accurate and detailed textures.
@@ 4.6 Ablation Study
Effectiveness of the Proposed Framework. We validate the effectiveness of the P2GAN structure as described in Table V for the denoising task. Here, V1 represents a GAN-based unsupervised denoising network with only unpaired synthesis. In addition, V2 extends V1 by adding the BGM loss, while V3 further includes the PL module. V4 introduces the branch U-S, and V5 is our baseline (P2GAN).
Table V shows that adding the BGM loss to the GAN-based unsupervised network (V1 to V2) results in a 0.17 dB PSNR improvement. Incorporating the PL module (V2 to V3) yields an additional PSNR increase of approximately 0.23 dB, highlighting the effectiveness of the PL module in enhancing synthetic image quality. When adding the “self-synthesis” constraint to V3 to obtain V4, there is a notable improvement of 0.45 dB PSNR on the SIDD Benchmark, indicating that combining “self-synthesis” with unpaired synthesis enhances network training and restorer performance. Comparing P2GAN with V4, we observe a substantial performance boost in P2GAN, with PSNR gains of 0.53 dB on the SIDD Validation. This demonstrates that P2GAN produces more realistic synthetic degraded images, improving the restorer’s performance.
@@
Effectiveness of the Number of Augmentation Images (N) in the Re-boosting (ReB) Module. Table VI shows that Reb-SC improves performance across all datasets. For $N$=2, the improvement is approximately 0.2 dB; for $N$=4, it is about 0.3 dB. However, when $N$=8, the performance improvement is less significant than $N$=4.